Calculate the number of equivalents in each of the following:

A.) 5 moles of K+

B.) 2 moles of OH−

C.) 5 moles of Ca2+

D.) 4 moles of CO2−3
Answer #1

A)

Magnitude of Charge = 1

Use:

Number of equivalent = number of mol * Magnitude of Charge

= 5*1

= 5

Answer: 5

B)

Magnitude of Charge = 1

Use:

Number of equivalent = number of mol * Magnitude of Charge

= 2*1

= 2

Answer: 2

C)

Magnitude of Charge = 2

Use:

Number of equivalent = number of mol * Magnitude of Charge

= 5*2

= 10

Answer: 10

D)

Magnitude of Charge = 2

Use:

Number of equivalent = number of mol * Magnitude of Charge

= 4*2

= 8

Answer: 8